# Is Mamba Capable of In-Context Learning?

### v1

@@ 3 Investigation of Simple NLP Tasks
In this section, we evaluated the ICL performance of various pre-trained Mamba language models, with parameter counts from 130 million to 2.8 billion.
@@
The pre-training for all variants was done on the Pile dataset (Gao et al., 2020), while the Mamba 2.8B (SP) checkpoint underwent further fine-tuning on ca. 600 billion tokens from the SlimPajama dataset.
We compared the Mamba variants to another RNN model with linear state dynamics named RWKV (Peng et al., 2023), which was also pre-trained on Pile [^1] We used RWKV v4 checkpoints trained on the Pile dataset by the authors available on huggingface., and popular transformer-based language models, such as LLama (Touvron et al., 2023), Pythia (Biderman et al., 2023), and GPT-J 6B (Wang & Komatsuzaki, 2021). We did not compare to S4 because we are not aware of S4 models pretrained on Pile.
We followed the experimental protocol of Hendel et al. (2023), which tested 27 NLP tasks spanning a wide range of categories, including algorithmic tasks (e.g., list element extraction), translation (e.g., English to Spanish), linguistic tasks (e.g., singular to plural conversion), and knowledge-based tasks (e.g., identifying country-capital pairs). For evaluation, we used the same datasets as Hendel et al. (2023) except for the algorithmic tasks, which were randomly generated (we use the same generation parameters). We report the mean accuracy over 400 generated test sets per task, each having five in-context examples.
@@
Figure 3 shows that ICL performance improves for all models with increasing number of parameters. Notably, Mamba 2.8B achieves a ICL accuracy close to LLama 7B, and on par with GPT-J and Pythia models. In addition, we find that Mamba consistently outperforms the similarly scalable RWKV at comparable parameter sizes. We provide a detailed table of per-task accuracies in Appendix A.3.
@@
Finally, we find that Mamba scales well with the number of in-context examples; see Figure 4. Particularly, Mamba 0.37B and 2.8B maintain a considerable performance edge over RWKV 0.43B.
4 Conclusions
@@
In this work, we have demonstrated that the recently proposed Mamba architecture is capable of effective ICL across tasks involving simple function approximation as well as more complex natural language processing problems. Our analysis showed that Mamba performs on par with transformer models, while also outperforming the S4 and RWKV baselines. We provide initial evidence that Mamba appears to solve ICL problems by incrementally refining its internal representations in a manner akin to an iterative optimization strategy, as transformer do. Overall, our findings establish Mamba as an efficient and performant alternative to transformers for ICL involving longer input sequences.
5 Acknowledgements
@@
This work was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828, and the Bundesministerium für Umwelt, Naturschutz, nukleare Sicherheit und Verbraucherschutz (BMUV, German Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection) based on a resolution of the German Bundestag (67KI2029A)
@@ Appendix A Appendix
A.1 Mamba Architecture
The Mamba architecture (Gu & Dao, 2023) is a sequence model that uses Structured State Space Models (SSMs) as its core building block. SSMs are recurrent neural networks with an high-dimensional latent state and linear state dynamics that feature also a “convolutional mode” which allows to parallelize the training training in a way similar to transformer models. Compared to previous SSMs, Mamba primary innovation is the introduction of a selection mechanism that makes the SSM parameters functions of the input, enabling selective propagation of information.
However, this selection mechanism poses a challenge for efficient convolutions. To address this, Mamba leverages a hardware-aware recurrent algorithm that exploits the GPU memory hierarchy, avoiding the need to materialize the full latent states in the GPU high bandwidth memory. Mamba’s architecture is designed to balance the trade-off between efficiency and effectiveness in sequence models by effectively compressing the input context into the latent state through selectivity.
Overall, the latent state in Mamba is updated as follows:
$h_{t}$ $=\bar{A}(\Delta(x_{t}))h_{t-1}+\bar{B}(x_{t},\Delta(x_{t}))x_{t}\quad,$
$y_{t}$ $=\bar{C}(x_{t})h_{t}\quad,$
where $h_{t}$, $x_{t}$ and $y_{t}$ are the current hidden states, current input and output respectively $\Delta,\bar{A},\bar{B},\bar{C}$ are specific parametric functions learned through backpropagation. See Gu & Dao (2023) for further details on Mamba’s architecture.
Mamba is referred to as linear time variant because $h_{t}$ is a linear function of $h_{t-1}$, but the linear transformation depends on the current input $x_{t}$. In contrast, previous SSM models, such as S4 and H3, were linear time invariant, since $\bar{A}$, $\bar{B}$ and $\bar{C}$ are not dependent on the current input.
A.2 ICL for simple function classes
@@
We follow the experimental setup of Garg et al. (2022) by building on their MIT-Licensed code at https://github.com/dtsip/in-context-learning.
A.2.1 Model parameters
@@
As in Garg et al. (2022), we used a GPT-2 (Radford et al., 2019) model with embedding size 256, 12 layers and 8 heads, resulting in 9.5 million parameters. As mentioned in the main text, we removed the positional embedding to improve the input length extrapolation. We set Mamba’s $D$ parameter to $256$, matching the transformer’s embedding dimension, while we set the number of Mamba’s layers to $24$, doubling the ones of the transformer ($12$). This is done because each Mamba block can be roughly seen as the fusion of a MLP and SSM, and has roughly half the number of parameters of a transformer block, which can be divided into an MLP and an attention part. For a fair comparison, S4 also uses $24$ layers, however we set the embedding size to 435 to match the 10 million parameters of Mamba and the transformer. We note that a similar comparison between S4 and transformer models, also testing for input length extrapolation, was done by  Lee et al. (2023). However, we used models with substantially more parameters (10 million vs. 500 thousand) and transformers without positional encoders.
A.2.2 Training details
@@
We adopted the same experimental setup used for the transformer model by (Garg et al., 2022) for Mamba, the transformer, and S4. At each training step, we computed the loss on a mini-batch of 64 prompts, each corresponding to a task sampled from a task distribution. We used no dropout in our experiments. We adopt a curriculum learning strategy: starting with training points of the lower-dimensional subspace and fewer input examples per prompt, and increasing the dimensionality of the subspace and the number of in-context examples every 2000 steps. For more details on the training, we refer to (Garg et al., 2022, A.2 Training).
For Mamba, we used the same learning rate used by Garg et al. (2022) for the transformer model (which is the only one we tried), which means that both Mamba and the transformer model undergo exactly the same training conditions. For S4 instead, we searched for the optimal learning rate in [1e-5, 1e-4, 1e-3, 1e-2] and selected 1e-3, which we found to achieve the lowest average loss on the query point at the end of training.
A.2.3 Simple task distributions
@@
Below, we describe how tasks are sampled for each task distribution that we considered. Task distributions are the same ones as in Garg et al. (2022).
@@
For each task we first sampled $x_{1},\dots,x_{k},x_{k+1}$ input points Then, for each point we sampled the output $y_{i}$ as a function of $x_{i}\in\mathbb{R}^{d}$, possibly adding noise. Finally, the prompt $P=x_{1},y_{1},\dots,x_{k},y_{k},x_{k+1}$ was passed as input to the model, which can be divided in context $x_{1},y_{1}),\dots,x_{k},y_{k}$, and query point $x_{k+1}$. We now describe how each input-output pair $(x_{i},y_{i})$ is computed for each task for different task distributions. For each task distribution we set $d=20$.
@@
Random quadrants
As linear regression, but the sign of each coordinate is randomly sampled, where every in-context example lies in one quadrant, while the query input lies in another.
@@
These networks represents functions of the form $f(x)=\sum_{i=1}^{r}\alpha_{i}\sigma(w_{i}^{T}x)$, where $\alpha_{i}\in\mathbb{R}$, $w_{i}\in\mathbb{R}^{d}$ and $\sigma(\cdot)=\max(0,\cdot)$ is the ReLU activation function. To generate a random prompt $P=(x_{1},f(x_{1}),...,x_{k},f(x_{k}),x_{k+1})$, we sample prompt inputs $x_{i}$’s from $N(0,I_{d})$, along with network parameters $\alpha_{i}$’s and $w_{i}$’s from $N(0,2/r)$ and $N(0,I_{d})$ respectively. We set the input dimension $d$ to 20 and the number of the hidden neurons $r$ to 100.
@@
A.2.4 Other baselines
As in Garg et al. (2022), we compared also with task distribution specific baselines, which we will discuss below. We refer to (Garg et al., 2022, Appendix A.3) for more details.
@@ A.2.4 Other baselines
It learns a decision tree greedily using scikit-learn’s decision tree regressor (Chen & Guestrin, 2016) with default parameters and max_depth equal to 2.
@@
We use the XGBoost library (Chen & Guestrin, 2016) to learn an ensemble of 50 decision trees with maximum depth 4 and learning rate 0.1.
A.2.5 Additional results
We provide additional results on Mamba and transformers on non-skewed linear regression in Figure 4(a). We show its out-of-distribution performance in Figure 6. As pointed out in the main text, we find that both Mamba and the transformer perform worse out-of-distribution compared to when they are trained on skewed linear regression. Furthermore, Figures 4(b) and 4(c) show that the performance of Mamba and transformers are very similar when trained and tested on decision trees and ReLU neural network tasks.
(a) Linear Regression
(b) Decision Tree
(c) ReLU Neural Network
Figure 5: Mamba compares with transformers when trained and tested on the same simple regression task distributions.
@@ A.2.5 Additional results
(d) Random Quadrants
Figure 6: OOD Distribution tests for Mamba and Transformer trained on (non-skewed) linear regression data
A.2.6 ICL learning curves
Figure 7: Results for grid search over $\gamma$ for GD++. We optimized $\gamma$ in order to have optimal performance of GD++ at iteration 24 (the same number of iterations reported in Figure 1(a)). We picked a value of
In our experimental setup, we employed a total of 40 in-context examples and evaluated the model on 30 distinct query examples. Each query example was assessed through an independent forward pass, resulting in 30 separate predictions made by processing a batch of 41 samples (in-context plus query) within the model, a method analogous to that described by Garg et al. (2022). Throughout these forward passes, we captured the feature activations (embeddings) from the intermediate layers for analysis. We carried out this analysis for 64 different tasks with 70 total examples each (40 in-context and 30 query examples).
We now describe the probing strategy in detail. Consider the input prompt $(x_{1},y_{1},x_{2},y_{2},...,x_{k},y_{k},x_{k+1})$, let $z_{l,i}$ be the model internal representation of the input token $x_{i}$ at layer $l$, and denote by $\hat{y}_{i}=g(z_{L,i})$ the model prediction for $x_{i}$, where $g$ is the decoder and $L$ is the total number of layers. For each layer $l$ we first obtain intermediate scalar predictions $\tilde{y}_{l,i}=g(z_{l,i})$. Since $g$ is not meant to be used on intermediate representations, we adjust the prediction on the query point $x_{k+1}$ by computing $\hat{y}_{l,k+1}=a_{l}g(z_{l,k+1})+b_{l}$, where the scalar scale and shift parameters $a_{l}$ and $b_{l}$ are obtained by least squares on the input-output pairs $(\tilde{y}_{l,k-n+1},y_{k-n+1}),(\tilde{y}_{l,k},y_{k})$, with $n=10$ and $k=40$. We can then measure the squared error $(\hat{y}_{l,k+1}-y_{k+1})^{2}$ between the intermediate predictions and the target.
(a) Skewed LR

### v2

@@ 4  Investigation of Simple NLP Tasks
In this section, we evaluated the ICL performance of various pre-trained Mamba language models, with parameter counts from 130 million to 2.8 billion. The pre-training for all variants was done on the Pile dataset (Gao et al.,, 2020), while the Mamba 2.8B (SP) checkpoint underwent further fine-tuning on ca. 600 billion tokens from the SlimPajama dataset.
We compared the Mamba variants to another RNN model with linear state dynamics named RWKV (Peng et al.,, 2023), which was also pre-trained on the Pile [^1] We used RWKV v4 checkpoints trained on the Pile dataset by the authors available on huggingface., and popular transformer-based language models, such as LLama (Touvron et al.,, 2023), Pythia (Biderman et al.,, 2023), and GPT-J 6B (Wang and Komatsuzaki,, 2021). We did not compare to S4 because we are not aware of S4 models pretrained on the Pile.
We followed the experimental protocol of Hendel et al., (2023), which tested 27 NLP tasks spanning a wide range of categories, including algorithmic tasks (e.g., list element extraction), translation (e.g., English to Spanish), linguistic tasks (e.g., singular to plural conversion), and knowledge-based tasks (e.g., identifying country-capital pairs). For evaluation, we used the same datasets as Hendel et al., (2023) except for the algorithmic tasks, which were randomly generated (we use the same generation parameters). We report the mean accuracy over 400 generated test sets per task, each having five in-context examples.
@@
Figure 6 shows that the ICL performance improves for all models with increasing number of parameters. Notably, Mamba 2.8B achieves an ICL accuracy close to LLama 7B, and on par with GPT-J and Pythia models. In addition, we find that Mamba consistently outperforms the similarly scalable RWKV at comparable parameter sizes. We provide a detailed table of per-task accuracies in Appendix A.2.
@@
Finally, we find that Mamba scales well with the number of in-context examples; see Figure 7. Particularly, Mamba 0.37B and 2.8B maintain a considerable performance edge over RWKV 0.43B and 3B, respectively.
5  Related Work
In-context learning capabilities in transformer-based models
Brown et al., (2020) introduced GPT-3 and were the first to show the remarkable in-context learning (ICL) abilities of large language models. They showed that larger models tend to utilize more in-context information compared to smaller models. Other work by Garg et al., (2022) showed that even small transformers exhibit ICL abilities on linear modeling tasks. These small transformer-based models have been subject to in-depth analysis by recent work (Dai et al.,, 2022; Von Oswald et al.,, 2023; Ahn et al.,, 2023; Bai et al.,, 2023) that showed that they can implement mechanisms akin to gradient descent or more complex optimization algorithms in the forward pass.
On a different tangent, Olsson et al., (2022) found that a two-layer attention-only network can develop the so-called “induction heads” mechanism, which outputs the token succeeding a previous instance of the current token, precisely when its ICL performance increases. Chan et al., (2022) investigated properties of the data-distribution which lead to the emergence of ICL abilities, while Reddy, (2024) identified factors for the abrupt emergence of the induction heads. Another line of work (Hendel et al.,, 2023; Todd et al.,, 2024) showed that some intermediate output of the forward pass of transformer models, named “task vectors”, encodes most of the information for the in-context task. In particular, merging a task vector with another from a different in-context task allows the model to solve a combination of the two tasks.
In-context learning benchmarks
Garg et al., (2022) tested transformers abilities to learn task distributions from simple function classes, like linear regression or Decision trees. Hendel et al., (2023) tested pre-trained LLMs on a suite of NLP tasks such as English to French translation. Lastly, Akyürek et al., (2024) proposed to test NLP models’ ability to learn formal languages.
Sub-quadratic transformer alternatives
Transformer-based models have a forward pass with time complexity scaling quadratically with respect to the input sequence length. To reduce the amount of computation, previous work proposed to introduce sparsity into the attention layers (Child et al.,, 2019; Qiu et al.,, 2019; Beltagy et al.,, 2020), where the tokens only attend a subset of the other tokens. Other works by Wang et al., (2020) proposed a low-rank factorization of the attention mechanism. Others proposed linear attention, where the attention weights are computed using possibly normalized dot-products without the softmax (Katharopoulos et al.,, 2020; Choromanski et al.,, 2020; Kasai et al.,, 2021; Peng et al.,, 2021; Yang et al., 2023b, ). Intriguingly, causal linear attention can be re-formulated either as a linear RNN or as the ratio between two linear RNNs, and hence its time complexity scales linearly with the length of the input sequence.
Recently, state-space sequence models gained significant interest due their strong performance. They are broadly related to recurrent neural networks with ideas from classical state space models, such as the Kalman filter (Kalman,, 1960), particle filters (Gordon et al.,, 1993), or hidden Markov models (Baum et al.,, 1970; Rabiner,, 1989). There exist various state-space models: S4 (Gu et al., 2021a, ), DSS (Gupta et al.,, 2022), S5 (Smith et al.,, 2023), GSS (Mehta et al.,, 2023), H3 (Fu et al.,, 2023), Selective S4 (Wang et al.,, 2023), or RWKV (Peng et al.,, 2023). While they have been successfully applied to audio and vision (Goel et al.,, 2022; Nguyen et al.,, 2022), they lacked performance for text. The recently proposed Mamba (Gu and Dao,, 2023) exhibits great performance in large language modeling and similar scaling properties as state of the art transformer models.
6  Broader Impact Statement
In-context learning (ICL) is an emergent phenomenon that drives the generalization skills of large language models (LLMs). While the widespread use of a poorly-understood technology like ICL may have many potential negative societal impacts, our work contributes to mitigate this problem by helping to build a mechanistic understanding of ICL. This understanding is crucial to build better, more explainable and more trustworthy AI systems in the future. Since the favourable scaling of Mamba (Gu and Dao,, 2023) makes it a strong candidate for dealing with long context sizes, we do believe it to be particularly important to understand the mechanisms behind ICL in this architecture, and we thus focus on this aspect in our paper.
7  Limitations
Our investigation into the Mamba architecture’s in-context learning (ICL) capabilities shows promise as an efficient alternative to transformer models for processing long input sequences, yet it is not without limitations. The study’s focus on simple function approximation and natural language processing (NLP) tasks raises questions about the generalizability of our findings to other domains like image or audio analysis, suggesting the need for broader research. The linear probing method used to understand Mamba’s approach may oversimplify the model’s complex optimization processes, indicating that more sophisticated probing techniques could yield deeper insights. Additionally, our evaluation covered a limited range of model sizes and configurations, and a more comprehensive comparison is necessary to fully assess Mamba’s performance relative to extensively optimized transformer models. The scalability of Mamba with increasing in-context examples and its computational efficiency also remain underexplored, highlighting areas for future investigation to better understand its practical applicability and potential across a wider range of tasks and settings.
8  Conclusions and Future Directions
@@
In this work, we have demonstrated that the recently proposed Mamba architecture is capable of effective in-context-learning (ICL) across tasks involving simple function approximation as well as more complex natural language processing problems. Our analysis showed that Mamba performs on par with transformer models, while also outperforming the S4 and RWKV baselines. We provide initial evidence that Mamba appears to solve ICL problems by incrementally refining its internal representations in a manner akin to an iterative optimization strategy, as transformers do. Overall, our findings suggest that Mamba can be an efficient and performant alternative to transformers for ICL involving longer input sequences.
In future work, it would be interesting to replace the attention-based backbone of existing in-context learned AutoML algorithms, such as TabPFN (Hollmann et al.,, 2023), ForecastPFN (Dooley et al.,, 2024) or Optformer (Chen et al.,, 2022), enabling them to handle longer sequences.
Acknowledgements
@@
This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828, and the Bundesministerium für Umwelt, Naturschutz, nukleare Sicherheit und Verbraucherschutz (BMUV, German Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection) based on a resolution of the German Bundestag (67KI2029A)
@@ Appendix A Appendix
A.1  ICL for simple function classes
@@
We follow the experimental setup of Garg et al., (2022) by building on their MIT-Licensed code at https://github.com/dtsip/in-context-learning.
A.1.1  Model parameters
@@
As in Garg et al., (2022), we used a GPT-2 (Radford et al.,, 2019) model with embedding size 256, 12 layers and 8 heads, resulting in 9.5 million parameters. As mentioned in the main text, we removed the positional encoding to improve the input length extrapolation. Mamba’s $N$ parameter is set to its default value $16$, while we set Mamba’s $D$ parameter to $256$, matching the transformer’s embedding dimension, while we set the number of Mamba’s layers to $24$, doubling the ones of the transformer ($12$). This is done because each Mamba block can be roughly seen as the fusion of a MLP and SSM, and has roughly half the number of parameters of a transformer block, which can be divided into an MLP and an attention part. For a fair comparison, S4 also uses $24$ layers, however we set the embedding size to 435 to match the 10 million parameters of Mamba and the transformer. We note that a similar comparison between S4 and transformer models, also testing for input length extrapolation, was done by  Lee et al., (2023). However, we used models with substantially more parameters (10 millions vs. 500 thousands) and transformers without positional encoders.
A.1.2  Training details
@@
We adopted the same experimental setup used for the transformer model by (Garg et al.,, 2022) for Mamba, the transformer, and S4. At each training step, we computed the loss on a mini-batch of 64 prompts, each corresponding to a task sampled from a task distribution. We used no dropout in our experiments. We adopt a curriculum learning strategy: starting with training points of the lower-dimensional subspace and fewer input examples per prompt, and increasing the dimensionality of the subspace and the number of in-context examples every 2000 steps. For more details on the training, we refer to (Garg et al.,, 2022, A.2 Training).
Differently for Garg et al., (2022), we use cosine annealing (from PyTorch) for Mamba, Transformer and S4, as we observed that it consistently improved performance.
The training was done Nvidia RTX 2080 GPUs and each training had a duration from 20 hours (for skewed/linear/sparse regression) up to 24 hours (for ReLU neural network and Decision tree). The evaluation for up to 500 ICL examples was done an Nvidia A100 GPU and took 16 hours for skewed/linear/sparse regression and up to 24 hours for ReLU neural network and Decision tree.
A.1.3  Simple task distributions
@@
Below, we describe how tasks are sampled for each task distribution that we considered. Task distributions are the same ones as in Garg et al., (2022).
@@
For each task we first sampled $x_{1},\dots,x_{k},x_{k+1}$ input points Then, for each point we sampled the output $y_{i}$ as a function of $x_{i}\in\mathbb{R}$, possibly adding noise. Finally, the prompt $P=x_{1},y_{1},\dots,x_{k},y_{k},x_{k+1}$ was passed as input to the model, which can be divided in context $x_{1},y_{1}),\dots,x_{k},y_{k}$, and query point $x_{k+1}$. We now describe how each input-output pair $(x_{i},y_{i})$ is computed for each task for different task distributions. We set the number of input dimensions $d=20$ for all task distributions.
@@
Sparse linear regression
First sample $w\sim\mathcal{N}(0,I_{d})$, then set all coordinates of $w$ except $s=3$ to zero. Finally, for $i=1,\dots,k+1$ sample $x_{i}\sim\mathcal{N}(0,I)$ and $y_{i}=w^{T}x_{i}$.
@@
Different orthants
As linear regression, but the sign of each coordinate is randomly sampled, where every in-context example lies in one quadrant, while the query input lies in another with high probability.
$d/2$-dimensional subspace
As linear regression, but the coordinates from $\lceil d/2\rceil$ to $d$ are set to 00 for each $x_{i}$.
@@
These networks represents functions of the form $f(x)=\sum_{i=1}^{r}\alpha_{i}\sigma(w_{i}^{T}x)$, where $\alpha_{i}\in\mathbb{R}$, $w_{i}\in\mathbb{R}$ and $\sigma(\cdot)=\max(0,\cdot)$ is the ReLU activation function. To generate a random prompt $P=(x_{1},f(x_{1}),...,x_{k},f(x_{k}),x_{k+1})$, we sample prompt inputs $x_{i}$’s from $N(0,I_{d})$, along with network parameters $\alpha_{i}$’s and $w_{i}$’s from $N(0,2/r)$ and $N(0,I_{d})$ respectively. We set the input dimension $d$ to 20 and the number of the hidden neurons $r$ to 100.
@@
A.1.4  Other baselines
As in Garg et al., (2022), we compared also with task distribution specific baselines, which we will discuss below. We refer to (Garg et al.,, 2022, Appendix A.3) for more details.
@@ A.1.4  Other baselines
Lasso
Fits a LASSO estimator to the in-context examples with a specified L1 regularization parameter.
@@
It learns a Decision tree greedily using scikit-learn’s Decision tree regressor (Chen and Guestrin,, 2016) with default parameters and max_depth equal to 2.
@@
We use the XGBoost library (Chen and Guestrin,, 2016) to learn an ensemble of 50 Decision trees with maximum depth 4 and learning rate 0.1.
A.1.5  Additional results
We provide additional results on Mamba and transformers on non-skewed linear regression in Figure 8. We show its out-of-distribution performance in Figure 9. Differently from the results in the main text, instead of cosine annealing we use the same learning rates used by Garg et al., (2022) for both Mamba and the transformer model. As pointed out in the main text, we find that both Mamba and the transformer perform worse out-of-distribution compared to when they are trained on skewed linear regression.
In Figure 10, we observe how the context length extrapolation performance is affected by using a higher number of examples during training for skewed linear regression tasks. In particular, the window where both models obtain good performance is wider the higher the number of in-context examples during training, with the transformer generally exhibiting a much lower degree of degradation.
Figure 8: Mamba compares with transformers when trained and tested on the same unskewed linear regression task distribution. We report one training run for each plot and method
@@ A.1.5  Additional results
(d) Different orthants
Figure 9: OOD Distribution tests for Mamba and Transformer trained on (non-skewed) linear regression data. We report one training run for each method.
(a) 40 examples
(b) 80 examples
(c) 160 examples ————————
Figure 10: Length extrapolation performance for varying number of in-context examples used for testing and training (dashed vertical lines) on the skewed linear regression task. We report 3 training runs for each plot and model
A.1.6  ICL learning curves
Figure 11: Results for grid search over $\gamma$ for GD++. We optimized $\gamma$ in order to have optimal average performance across tasks of GD++ at iteration 24 (the same number of iterations reported in Figure 3(d)). We picked a value of approximately $2\times 10^{-4}$
(a) Sparse LR